# Kernel Three Pass Regression Filter

### v3

@@ 5 Empirical Applications
We apply our proposed method to real-world applications, focusing on forecasting time series variables across various economic domains such as national income, finance, labor, housing, prices, etc. To assess the performance of our approach, we conduct comparative analyses against competitive methods, employing the out-of-sample $R^{2}$ performance metric as a benchmark. Out of sample $R^{2}$ is computed as:
@@
It computes the out-of-sample proximity of our forecast $\hat{y}$ with the target ($y$) relative to a historical mean $(\bar{y})$; a positive value indicates that the forecast is better than the historical mean. Detailed explanations of performance metrics computation are provided in the Supplementary appendix-B.2.
We compare our method against six different forecasting methods. The first is the PC regression proposed by Stock & Watson, (2002b); which we write as PC in our performance tables, PC-Squared (PC-Sq) and Squared-PC(Sq-PC) of Bai & Ng, (2008), kernel PCA (kPCA) [Kutateladze, (2022)], our linear counterpart, the 3PRF, and autoregressive model of lag order two [^8] We compared the performance across various autoregressive lags and found minimal differences. However, a lag of two generally showed superior performance in most cases.. Some of these methods require tuning of hyper-parameters to provide the best results, we do tune them as discussed in the subsection-5.1.2.
As discussed in section 2, different kernel functions correspond to different $\varphi(\cdot)$. We use the radial basis function (RBF) kernel.
@@
where $\sigma$ is a hyperparameter determined via cross-validation. The use of this kernel is justified by its strong performance in macroeconomic forecasting, as documented by rbf_ref1, Exterkate et al., (2016), and Kutateladze, (2022).
@@ 5.1 Data and Hyper-Parameter Tuning
We utilize the quarterly macroeconomic dataset, FRED-QD. It spans the time period 1959-2023. This dataset encompasses a comprehensive array of more than 250 variables, including macroeconomic (such as GDP, Consumption, and Investment), financial, labor market, housing, and industrial and manufacturing variables. We present a tabulation comprising the mnemonic codes and details of the variables in the FRED-QD dataset alongside their counterparts in the Stock-Watson dataset in Supplementary appendix-B.4 for the series we forecast in this section.
@@ 5.1.1 Data Transformation
Estimation exercises involving nonstationary time series pose significant challenges. Nonstationarity is ubiquitous in economic and financial data. Nonstationary variables lack a defined population mean, and the sample standard deviation tends to diverge as the number of observations increases, see Onatski & Wang, (2021) for a more detailed discussion. Generally, researchers address this by manually examining each series to identify necessary transformations before computing principal components. Hamilton & Xi, (2024) offers an improved method for transforming the predictors to achieve stationarity. We use their method to make our data stationarity.
Scholars in the literature often employ sample periods devoid of structural breaks. Fan et al., (2023) notes that “There exist significant structural breaks for many variables around the year of the financial crisis in 2008 which makes our data non-stationary even after performing the suggested transformations”. Therefore, our study focuses on the stationary period spanning from 1965 to 2007999Another indirect advantage of the choice of this sample period is that it gives us the number of samples less than the number of predictors ($T<N$), hence a truly high-dimensional scenario to test our method in.. We conduct analyses on different combinations of the sample periods, including the entire available sample period from 1959 to 2023, and find no qualitative discrepancies in our findings.
@@ 5.1.2 Hyper-parameter Tuning
Our methodology incorporates the kernel as a fundamental element of the estimation process. The kernel function includes a hyperparameter that necessitates optimization. Concurrently, a similar hyperparameter requires tuning in the context of a competitor method, namely kernel PCA. Thus, we employ an identical tuning procedure for both methodologies. We adopt a RBF kernel, which relies on a single hyperparameter, denoted as $\sigma$, for our specific applications. We partition the data into two folds and conduct cross-validation to determine the optimal tuning parameter. Further elaboration on the algorithm employed for this purpose is provided in the Supplementary appendix-B.5.
Furthermore, among our competitive methodologies, where factors are computed as PCs, we are required to specify the number of factors. To address this, we employ the eigenvalue ratio test method proposed by Ahn & Horenstein, (2013). This method computes the ratio of each eigenvalue to its predecessor and selects the number of principal components corresponding to the index where this ratio attains its maximum value. We employ a single factor throughout our analyses in both the 3PRF and kernel 3PRF models. This choice is often prudent within the 3PRF setting, as elucidated by Kelly & Pruitt, (2015), who highlight instances where a single factor can effectively represent a multi-factor system. When factors exhibit the same variances [^10] See appendix section A.7.2 in Kelly & Pruitt, (2015), a single proxy achieves optimal performance, and even when variances are not identical but closely aligned, one factor estimated through a single auto proxy typically explains a significant portion of the variation [^11] See simulations in appendix A.7.3 in Kelly & Pruitt, (2015), rendering residual variation minimal. While we assessed the performance of our estimator with varying numbers of factors, we consistently observed that a single factor predominates, thus, we report results based on this configuration.
@@ 5.2 Forecasting Using Theory Guided Proxies
The primary objective of this subsection is to establish the viability of theory-guided proxies in forecasting using our method. We also compare with the linear benchmark, i.e., 3PRF of Kelly & Pruitt, (2015). A more extensive performance evaluation will be presented in subsequent subsections, where we used the auto-proxy method discussed in table-3 to construct forecasts using Kernel 3PRF.
@@ 5.2.1 Forecasting GDP Using Investment and Consumption
We construct GDP forecasts using Consumption and investment as proxies and report the results in Table-4.
@@
This exercise proves the efficacy of K3PRF over 3PRF while employing theory-guided proxies. Furthermore, our method outperforms the nearest competitive method Kelly & Pruitt, (2015).
@@ 5.2.2 Forecasting Inflation using Quantity Theory of Money
We reproduce the theory-guided proxy example discussed in Kelly & Pruitt, (2015). Δ(Price $\Delta(\text{Price level})$ i.e. inflation is our target variable for forecasting. The results for one-period ahead inflation forecasts are presented in Table-5. The quantity theory of money equation states that:
Δ(Money supply)×Δ(Velocity of money)Δ(Real Product)=Δ(Price $\frac{\Delta(\text{Money supply})\times\Delta(\text{Velocity of money})}{\Delta(\text{Real Product})}=\Delta(\text{Price level})$
@@
The results indicate that the theory-guided proxies effectively capture inflation dynamics, yielding performance comparable to that of the closest competitor. It is important to emphasize again that this analysis focuses on one-step-ahead forecasts, which are not the primary strength of our methodology. The purpose of presenting these results is solely to demonstrate the workings of the procedure through the theory-guided proxies.
@@ 5.3 Comparative Forecasting Plots
To visually demonstrate the enhanced performance of kernel 3PRF compared to its linear counterpart, we provide comparative performance plots across four distinct types of economic series spanning various domains: macroeconomic series (Exports), price series (GDP Deflator), manufacturing series (Industrial Production), and financial series (S&P 500 Index) in figure-3 and 4. Plots of all other series on different forecast horizons are given in the Supplementary appendix-B.6.
@@ 5.4 Forecasting Aggregate Macroeconomic Variables
An astute economic decision, such as monetary policy formulation, hinges upon well-informed anticipations of future trends in macroeconomic and financial data. Consequently, forecasting macroeconomic variables emerges as a pivotal pursuit for economists. Quoting Federal Reserve of New York’s website, Kim & Swanson, (2014) notes, “In formulating the nation’s monetary policy, the Federal Reserve considers a number of factors, including the economic and financial indicators which follow, as well as the anecdotal reports compiled in the Beige Book. Real Gross Domestic Product (GDP); Consumer Price Index (CPI); Nonfarm Payroll Employment Housing Starts; Industrial Production/Capacity Utilization; Retail Sales; Business Sales and Inventories; Advance Durable Goods Shipments, New Orders and Unfilled Orders; Lightweight Vehicle Sales; Yield on 10-year Treasury Bond; S&P 500 Stock Index; M2”. We, therefore, aim to forecast some of these crucial indicators in this paper. We compare the performance of our model against the competitors. This section forecasts seven macro series: GDP, Consumption, Investment, Exports, Imports, Fixed Investment, and Industrial Production (Final).
@@
To present the results in an organized manner, we create two tables. In Table-6, we display the forecasting performance for three series: GDP, Consumption, and Investment, which we informally refer to as ‘Group-I’. Table-7 presents a comparative analysis of forecasting performance for ‘Group-II’ macro variables121212 Variables’ FRED-QD code and description can be found in Supplementary appendix-B.4: Exports, Imports, Fixed Investments, and Industrial Production (Final Index). As defined earlier in the text, the reported numbers in the tables represent out-of-sample $R^{2}$ values across various forecast horizons ranging from one period ahead to twelve periods ahead.
Results highlight a secular observation that among various unsupervised forecasting methodologies, PC, Squared-PC, PC-squared, and non-linear unsupervised approaches such as kernel PCA, none exhibit superior performance compared to our proposed method across any forecast horizon for the seven series under consideration. While the supervised linear forecasting model 3PRF demonstrates improved performance relative to the unsupervised techniques, it still falls short of outperforming our non-linear supervised approach. Notably, the autoregressive (AR) model emerges as the sole contender capable of surpassing our method in the shorter horizons, albeit only marginally and for a few series. However, our method significantly outperforms all the competitors across longer horizons. Therefore, our method emerges as a dependable and preferred forecasting framework across all forecast horizons in macroeconomic prediction tasks.
@@ 5.5 Forecasting Labor Market and Price Variables
Our relative forecast performance results are qualitatively similar to those of aggregate macroeconomic series forecasting.
@@ 5.6 Forecasting Housing and Financial Variables
As seen from Table-9, forecasting HStart(total) proves to be a difficult problem. While most forecasting methods do not beat the historical average, our method performs better than all other methods at all horizons. It is relatively easy to forecast housing in the western census region, and our method performs better than all other methods except for a few cases. We find similar patterns in financial market variables, thereby omitting discussion.
@@ 6.1 Description of Comparisons
Our investigation encompasses the comparative performance of models across a total of $176\times 8=1408$ target-horizon combinations. The results of these comparisons, indicating the percentage of instances where a particular method demonstrated superior performance, are presented in Supplementary appendix-B.7. For example, if a method emerged as the best performer in 704 out of 1408 combinations, it would be represented by a value of 50 in the table. Essentially, we list the relative frequency of the occurrence of the best performance of a given method.
@@
The “All Horizons” set of rows summarizes all 1408 comparisons, i.e. encompassing all horizons and all series. Recognizing that forecast objectives may vary in time horizon, we scrutinize comparative performances in short- and long-run contexts. The “Short-run” rows incorporate horizons $h={1,2,3,4}$, comprising 708 (calculated as $176\times 4$) combinations, while the “Long-run” row includes horizons $h={6,8,10,12}$, similarly amounting to 708 combinations. Additionally, the portion labeled as “Excluding AR” excludes the auto-regressive method and compares the remaining methods across all 1408 combinations. For more granular analysis, we report comparative performance numbers for each forecast horizon $h$. These numbers are reported in the Supplementary appendix-B.7.
It is important to note that multiple ‘best’ methods may exist for a non-zero tolerance level, resulting in the sums of rows (in Table 5 in Supplementary appendix-B.7) exceeding 100 percent. However, for a tolerance level of zero, the rows sum to 100 percent.

### v4

@@ 5 Empirical Applications
We apply our proposed method to real-world applications, focusing on forecasting time series variables across various economic domains such as national income, finance, labor, housing, prices, etc. To assess the performance of our approach, we conduct comparative analyses against competitive methods, employing the out-of-sample $R^{2}$ performance metric as a benchmark. The out-of-sample $R^{2}$ metric is computed as:
@@
It computes the out-of-sample proximity of our forecast $\hat{y}$ with the target ($y$) relative to a historical mean $(\bar{y})$; a positive value indicates that the forecast is better than the historical mean. Detailed explanations of performance metrics computation are provided in the Online Appendix-B.2.
5.1 Overview
This section provides an introduction to our competing methods and then a high-level summary of the empirical applications.
5.1.1 Competing Methods
We compare our method against eight different forecasting approaches. The first is the principal components (PC) regression proposed by stock2002diff, denoted as PC in our performance tables. We also consider PC-Squared (PC-Sq) and Squared-PC (Sq-PC) from bai2008, kernel PCA (kPCA) [kpca], our linear counterpart 3PRF (kelly), and the autoregressive (AR) model with optimal lag pp. [^8] We evaluate several choices of $p$ and report the best-performing one. This represents the ideal performance for AR models in a setting where the true lag is known; in practice, $p$ cannot be chosen based on maximum out-of-sample $R^{2}$.
The Diffusion Index (DI) serves as a benchmark forecasting method that combines AR and PCA. We select the number of principal components using the eigenvalue ratio test of ahn2013eigenvalue and run several DIs for different AR($p$), retaining the one with the $p$ that achieves the highest out-of-sample $R^{2}$. In this sense, we assess the upper bound of DI performance, since in practice the optimal $p$ cannot be chosen based on out-of-sample results.
Finally, we consider the Autoencoder (AE), a modern neural network method that learns a latent-space or manifold representation (AE_hinton). The crucial hyperparameter, the number of latent dimensions, is chosen via cross-validation, consistent with the hyperparameter tuning applied to the other methods. AE is conceptually the closest neural-network-based competitor to our approach, as it also learns a latent representation of the data. Some of these methods require hyperparameter tuning to achieve optimal results; we perform such tuning as described in subsection 5.2.2.
As discussed in Section 2, different kernel functions correspond to different $\varphi(\cdot)$. We use the Gaussian [^9] The Gaussian kernel is mathematically equivalent to the commonly used RBF kernel in some papers. kernel throughout our applications.
@@ 5.1.1 Competing Methods
where $\sigma$ is a hyperparameter determined via cross-validation. The use of this kernel is justified by its strong performance in macroeconomic forecasting, as documented by rbf_ref1, k_ridge, and Varlam.
5.1.2 Summary of Empirical Applications
We begin by discussing the data sources, data transformations for stationary series, and the hyperparameter tuning procedure. The first empirical application illustrates the working of our method using theory-guided proxies. We then provide visual evidence showing how our method captures the nonlinearities present in the target variable compared with its nearest competitor. Next, we present the out-of-sample forecasting performance for several key variables of interest across a broad range of economic categories. Finally, we compare our method against others in a comprehensive forecasting exercise using all available series in the dataset as a target.
@@ 5.2 Data and Hyper-Parameter Tuning
We utilize the quarterly macroeconomic dataset, FRED-QD. It spans the time period 1959-2023. This dataset encompasses a comprehensive array of more than 250 variables [^10] There are 176 series for which we can find no missing values from 1959 to 2023, therefore our number of time series used are 176 in this analysis., including macroeconomic (such as GDP, Consumption, and Investment), financial, labor market, housing, and industrial and manufacturing variables. We present a tabulation comprising the mnemonic codes and details of the variables in the FRED-QD dataset alongside their counterparts in the Stock-Watson dataset in Online Appendix-B.4 for the series we forecast in this section.
@@ 5.2.1 Data Transformation
Estimation exercises involving nonstationary time series pose significant challenges. Nonstationarity is ubiquitous in economic and financial data. Nonstationary variables lack a defined population mean, and the sample standard deviation tends to diverge as the number of observations increases, see Onatski_Wang for a more detailed discussion. Generally, researchers address this by manually examining each series to identify necessary transformations before computing principal components. hamilton_tranformation offers an improved method for transforming the predictors to achieve stationarity. We use their method to make our data stationarity.
Scholars in the literature often employ sample periods devoid of structural breaks. fan2023latent notes that “There exist significant structural breaks for many variables around the year of financial crisis in 2008 which makes our data non-stationary even after performing the suggested transformations”. Therefore, our study focuses on the stationary period spanning from 1965 to 2007111111Another indirect advantage of the choice of this sample period is that it gives us the number of samples less than the number of predictors ($T<N$), hence a truly high-dimensional scenario to test our method in.. For robustness checks, we examine three other sample periods. First, 1965–2019, which captures the Global Financial Crisis (GFC), allowing us to assess performance during high-volatility periods, as the GFC falls within the test sample under our 70% training split. Second, 1965–2023, the full sample [^12] Although our data begins in 1959, we exclude early observations to accommodate AR(p) models with sufficiently large lags., encompassing both high- and low-volatility periods, including the GFC, COVID-19, and the Great Moderation. Third, 1984–2007, which corresponds to the Great Moderation and enables the evaluation of performance in a low-volatility environment. We found that the results are qualitatively similar across all sub-sample periods and convey the same story as in the main text (The results are tabulated in Online Appendix B.9). These findings indicate that our method is robust to both high-volatility (turbulent) and low-volatility (Great Moderation) periods.
@@ 5.2.2 Hyper-parameter Tuning
Our methodology incorporates the kernel as a fundamental element of the estimation process. The kernel function includes a hyperparameter that necessitates optimization. Concurrently, a similar hyperparameter requires tuning in the context of a competitor method, namely kernel PCA. Thus, we employ an identical tuning procedure for both methodologies. We use Gaussian kernel, which relies on a single hyperparameter, denoted as $\sigma$, for our specific applications. We partition the data into two folds and conduct cross-validation to determine the optimal tuning parameter. Further elaboration on the algorithm employed for this purpose is provided in the Online Appendix-B.5. We need to choose an optimal lag $p$ for the autoregressive (AR) and Diffusion Index (DI) methods, which we have already discussed. For the Autoencoder (AE) method, the key tuning parameters include the number of layers, layer size, and the dimension of the latent space, of which the latent space dimension is the most critical. We conduct experiments by varying the number of layers and layer sizes, and determine the optimal latent space dimension through cross-validation. Furthermore, a similar cross-validation procedure is applied to determine other hyperparameters in competing methods to ensure a fair comparison.
Among our competitive methodologies, where factors are computed as PCs, we are required to specify the number of factors. To address this, we employ the eigenvalue ratio test method proposed by ahn2013eigenvalue. This method computes the ratio of each eigenvalue to its predecessor and selects the number of principal components corresponding to the index where this ratio attains its maximum value. We employ a single factor throughout our analyses in both the 3PRF and kernel 3PRF models. This choice is often prudent within the 3PRF setting, as elucidated by kelly, who highlight instances where a single factor can effectively represent a multi-factor system. When factors exhibit the same variances [^13] See appendix section A.7.2 in kelly, a single proxy achieves optimal performance, and even when variances are not identical but closely aligned, one factor estimated through a single auto proxy typically explains a significant portion of the variation [^14] See simulations in appendix A.7.3 in kelly, rendering residual variation minimal. While we assessed the performance of our estimator with varying numbers of factors, we consistently observed that a single factor predominates; thus, we report results based on this configuration.
@@ 5.3 Forecasting Using Theory Guided Proxies
The primary objective of this subsection is to establish the viability of theory-guided proxies in forecasting using our method. We also compare with the linear benchmark, i.e., 3PRF of kelly. A more extensive performance evaluation will be presented in subsequent subsections, where we used the auto-proxy method discussed in table-2 to construct forecasts using Kernel 3PRF.
@@ 5.3.1 Forecasting GDP Using Investment and Consumption
Consumption and investment are well-known drivers of GDP, and they likely capture the underlying latent factors that drive future economic activity. Using these variables as theory-guided proxies allows us to extract the relevant predictive components from the broader cross section of macroeconomic predictors. We construct GDP forecasts using these variables as proxies and report the results in Table-3.
@@
This exercise proves the efficacy of theory-guided proxies. Our method outperforms the nearest competitive method kelly.
@@ 5.3.2 Forecasting Inflation using Quantity Theory of Money
[…]
$\frac{\Delta(\text{Money supply})\times\Delta(\text{Velocity of money})}{\Delta(\text{Real Product})}=\Delta(\text{Price level})$
[…]
@@
The results indicate that the theory-guided proxies successfully capture the key inflation dynamics, achieving performance comparable to the linear benchmark 3PRF. Importantly, this also demonstrates that in settings where the relationship is approximately linear, the kernel 3PRF can adapt and produce forecasts of similar quality, as appropriate tuning of the hyperparameters allows it to effectively capture linear effects. It is important to emphasize again that this analysis focuses on one-step-ahead forecasts, which are not the primary strength of our methodology. The purpose of presenting these results is solely to demonstrate the workings of the procedure through the theory-guided proxies.
@@ 5.4 Comparative Forecasting Plots
To visually demonstrate the enhanced performance of kernel 3PRF compared to its linear counterpart, we provide comparative performance plots across four distinct types of economic series spanning various domains: macroeconomic series (Exports), price series (GDP Deflator), manufacturing series (Industrial Production), and financial series (S&P 500 Index) in figure-3 and 4. Plots of all other series on different forecast horizons are given in the Online Appendix-B.6.
@@ 5.5 Forecasting Aggregate Macroeconomic Variables
An astute economic decision, such as monetary policy formulation, hinges upon well-informed anticipations of future trends in macroeconomic and financial data. Consequently, forecasting macroeconomic variables emerges as a pivotal pursuit for economists. Quoting Federal Reserve of New York’s website, kim_macro_forecasts notes, “In formulating the nation’s monetary policy, the Federal Reserve considers a number of factors, including the economic and financial indicators which follow, as well as the anecdotal reports compiled in the Beige Book. Real Gross Domestic Product (GDP); Consumer Price Index (CPI); Nonfarm Payroll Employment Housing Starts; Industrial Production/Capacity Utilization; Retail Sales; Business Sales and Inventories; Advance Durable Goods Shipments, New Orders and Unfilled Orders; Lightweight Vehicle Sales; Yield on 10-year Treasury Bond; S&P 500 Stock Index; M2”. We, therefore, aim to forecast some of these crucial indicators in this paper. We compare the performance of our model against the competitors. This section forecasts seven macro series: GDP, Consumption, Investment, Exports, Imports, Fixed Investment, and Industrial Production (Final).
@@
To present the results in an organized manner, we create two tables. In Table-5, we display the forecasting performance for three series: GDP, Consumption, and Investment, which we informally refer to as ‘Group-I’. Table-6 presents a comparative analysis of forecasting performance for ‘Group-II’ macro variables [^15] Variables’ FRED-QD code and description can be found in Online Appendix-B.4: Exports, Imports, Fixed Investments, and Industrial Production (Final Index). As defined earlier in the text, the reported numbers in the tables represent out-of-sample $R^{2}$ values across various forecast horizons ranging from one period ahead to twelve periods ahead.
Results highlight a secular observation that among various unsupervised forecasting methodologies—PC, Squared-PC, PC-Squared, and non-linear unsupervised approaches such as kernel PCA—none exhibit superior performance compared to our proposed method across any forecast horizon for the seven series under consideration. While the supervised linear forecasting model 3PRF demonstrates improved performance relative to the unsupervised techniques, it still falls short of our non-linear supervised approach. Notably, the autoregressive (AR) model and its variant, the Diffusion Index (DI), emerge as the sole contenders capable of surpassing our method at shorter horizons, albeit only marginally and for a few series. The neural-network-based competitor, the Autoencoder (AE), does not outperform our method at any forecast horizon—a result that is not surprising, as neural-network-based methods are known to require large sample sizes for effective training, which is not the case here.
In summary, our method significantly outperforms all competitors across longer horizons and competes well with AR(p) and DI in the short run. Therefore, it serves as a dependable and preferred forecasting framework across all forecast horizons in macroeconomic prediction tasks.
@@ 5.6 Forecasting Labor Market and Price Variables
While the major story remains of the forecast performance results are qualitatively similar to those of aggregate macroeconomic series forecasting. We want to highlight the following: in the Unemployment Rate and GDP Deflator, our method beats AR(p) and DI method even in the short-run, barring $h=1$, and in the remaining two variables, our method is the best except initial two horizons.
@@ 5.7 Forecasting Housing and Financial Variables
We observe that the linear AR(p) and DI models perform exceptionally well for housing variables, with our method closely trailing them while outperforming all other competitors. In contrast, for financial variables, our method decisively surpasses all competing approaches as the forecast horizon extends to $h\geq 3$.
5.8 Robustness Check: Evaluating Forecasts After Partialing Out AR(4) Variation
We observe that the AR(p) model and its closely related method, DI, perform particularly well in the short run. It is well-established in the literature that AR(p) models are difficult to outperform over shorter forecast horizons. Possibly for this reason, some forecasting approaches evaluate performance after accounting for the variation captured by AR(p). For instance, our closest competitor kelly effectively forecasts the innovation of an AR(4) process for each variable, thereby “partialing out” the predictive information contained in the target variable’s own lags.
Following a similar approach, we also conduct an exercise in which we partial out AR(4) variation and subsequently compare the forecasting performance of all methods. We find that the qualitative patterns remain consistent, and in some cases are even more pronounced. The detailed results are presented in Online Appendix B.8.
@@ 6.1 Description of Comparisons
Our investigation encompasses the comparative performance of models across a total of $176\times 8=1408$ target-horizon combinations. The results of these comparisons indicate the percentage of instances where a particular method demonstrated superior performance to all other competitors. For example, if a method emerged as the best performer in 704 out of 1408 combinations, it would be represented by a value of 50 in the table. Essentially, we list the relative frequency of the occurrence of the best performance of a given method.
@@
The “All Horizons” set of rows summarizes all 1408 comparisons, i.e., encompassing all horizons and all series. Recognizing that forecast objectives may vary in time horizon, we scrutinize comparative performances in short- and long-run contexts. The “Short-run” rows incorporate horizons h=1,2,3,4h={1,2,3,4}, comprising 708 (calculated as $176\times 4$) combinations, while the “Long-run” row includes horizons h=6,8,10,12h={6,8,10,12}, similarly amounting to 708 combinations. Additionally, the portion labeled as “Excluding AR” excludes the auto-regressive AR(p) and closely related Diffusion Index (DI) methods and compares the remaining methods across all 1408 combinations. For more granular analysis, we report comparative performance numbers for each forecast horizon $h$. These numbers are reported in the Online Appendix-B.7.
It is important to note that multiple ‘best’ methods may exist for a non-zero tolerance level, resulting in the sums of rows (in Table 5 in Online Appendix-B.7) exceeding 100 percent. However, for a tolerance level of zero, the rows sum to 100 percent.